# Interplay between local and global order controls bifurcations in nonreciprocal systems without a conservation law

### v1

@@ Abstract
We have shown that up to first order the static long-range order can be written as $m^{\mu}_{\rm s}=m^{\mu}_{0,\rm s}+K\delta m^{\mu}_{\rm s}+\mathcal{O}(K^{2})$ under a small perturbation in $K$, where $m^{\mu}_{0,\rm s}\neq 0$ for $J>\ln{(1+\sqrt{2})}/2$ as shown in Eq. (S97). Furthermore, the perturbation $\delta m^{\mu}_{\rm s}$ is bounded by a finite number independent of $K$, as shown in Eq. (S110). This means that for arbitrary small nonzero $K$, the perturbed long-range order gets arbitrary close to the unperturbed value, i.e.
$|m^{\mu}_{\rm s}-m^{\mu}_{0,\rm s}|\leq|K|\left(\frac{1+\ln{(c/2)}}{b}\right).$ (S111)
@@
If the static long-range order vanished for any arbitrary nonzero $K$ on the square lattice, it would indicate that $m^{\mu}_{\rm s}$ is a discontinuous and non-differentiable function of $K$ for $J>\ln{(1+\sqrt{2})}/2$, as it would suddenly jump from $m^{\mu}_{0,\rm s}\neq 0$ to $m^{\mu}_{\rm s}=0$ in the limit $K\rightarrow 0^{+}$. This cannot be true, since we have established that $m^{\mu}_{\rm s}$ gets arbitrarily close to $m^{\mu}_{0,\rm s}$ up to first order in $K$. This invalidates the statement that static long-range order is destroyed in the perfectly nonreciprocal Ising model for any $K\neq 0$.
@@
S7 Kinetic Monte-Carlo simulations
@@
Figure S2: Oscillations in the long-range order $m^{\mu}(t)$ on the mean-field lattice (a), Bethe lattice (b), and square lattice (c), obtained by Monte-Carlo simulations. In each panel, we plot 10 independent trajectories for the magnetization on lattice $a$ (green) and $b$ (blue). Fluctuations between the trajectories lead to small differences over time, which are decreasing with increasing system size, a signature of self-averaging. The values for $J$ and $K$ for each of the lattices are reported in table S1.
@@
In Table S1 we summarize the system size, number of trajectories, and parameter settings, which were used to obtain the spectral density shown in Fig. 3 in the Letter. As initial conditions, we selected a randomly mixed configuration of up and down spins for fixed magnetization. A subset of trajectories for the short-range order are displayed in Fig. S2, and in each of the lattices we observe coherent oscillations.
@@
S7.2 Eigenvalues of Laplacian matrix

### v2

@@ Abstract
We have shown that up to first order the static global order can be written as $m^{\mu}_{\rm s}=m^{\mu}_{0,\rm s}+K\delta m^{\mu}_{\rm s}+\mathcal{O}(K^{2})$ under a small perturbation in $K$, where $m^{\mu}_{0,\rm s}\neq 0$ for $J\gtrsim\ln{(1+\sqrt{2})}/2$ as shown in Eq. (S96). Furthermore, the perturbation $\delta m^{\mu}_{\rm s}$ is bounded by a finite number independent of $K$, as shown in Eq. (S109). This means that for arbitrarily small nonzero $K$, the perturbed global order gets arbitrary close to the unperturbed value, i.e.
$|m^{\mu}_{\rm s}-m^{\mu}_{0,\rm s}|\leq|K|\left(\frac{2N(1+\ln{(c/2)})}{b\tau}\right).$ (S110)
To be more precise, if we set
$0<|K|\leq\mathcal{O}(1/N),$ (S111)
@@
the perturbed steady-state magnetization must come arbitrarily close to the unperturbed steady-state magnetization by merely changing the proportionality factor. If the static global order vanished for any arbitrary nonzero $K$ on the square lattice, it would indicate that $m^{\mu}_{\rm s}$ is a discontinuous and non-differentiable function of $K$ for $J\gtrsim\ln{(1+\sqrt{2})}/2$, as it would suddenly jump from $m^{\mu}_{0,\rm s}\neq 0$ to $m^{\mu}_{\rm s}=0$ in the limit $K\rightarrow 0^{+}$. This cannot be true, since we have established that $m^{\mu}_{\rm s}$ gets arbitrarily close to $m^{\mu}_{0,\rm s}$ up to first order in $K$. Hence, the static global order is not destroyed for finite system sizes ($N<\infty$) in the perfectly nonreciprocal Ising model at least when $0<K\leq\mathcal{O}(1/N)$.
@@
S6.6 What happens in the thermodynamic limit?
What happens to the static global order when we take $N\rightarrow\infty$? It is known that the infinite volume Glauber dynamics is not ergodic as the Gibbs measure is not unique [7]. Therefore, one cannot use Theorem 1. However, in the presence of a magnetic field with $h>0$ there exists a region $J\geq J_{0}$ (this includes the infinite $J\rightarrow\infty$ limit) such that the infinite volume Glauber dynamics has exponential convergence, i.e., is weakly ergodic, as shown in Theorem 5.1b in [8] (page 479). Hence, in this regime we can use Theorem 1 to obtain a bound for $|K|\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$, which would also be given by Eq. (S108) (albeit that $b$ and $c$ might change due to the presence of $h>0$). Looking at Eq. (S110) we can then take a scaling limit $N\rightarrow\infty$ and $|K|N={\rm constant}$, such that the perturbed steady-state long-range order can get arbitrarily close to the unperturbed value by simply changing the proportionality constant. Hence, we conclude that there exists a regime for $|K|N={\rm constant}$ where the static order is preserved when
$N\xrightarrow[h>0,\ J\geq J_{0}]{|K|N={\rm const.}}\infty.$ (S112)
Appendix S7 Kinetic Monte-Carlo simulations
@@
Figure S2: Oscillations in the global order $m^{\mu}(t)$ on the mean-field lattice (a), Bethe lattice (b), and square lattice (c), obtained by Monte-Carlo simulations. In each panel, we plot 10 independent trajectories for the magnetization on lattice $a$ (green) and $b$ (blue). Fluctuations between the trajectories lead to small differences over time, which are decreasing with increasing system size, a signature of self-averaging. The values for $J$ and $K$ for each of the lattices are reported in table S1.
@@
In Table S1 we summarize the system size, number of trajectories, and parameter settings, which were used to obtain the spectral density shown in Fig. 3 in the Letter. As initial conditions, we selected a randomly mixed configuration of up and down spins for fixed magnetization. A subset of trajectories for the global order are displayed in Fig. S2, and in each of the lattices we observe coherent oscillations.
@@
S7.2 Videos
To visualize the dynamics of the nonreciprocal Ising model on the mean-field lattice, Bethe lattice, and square lattice, we have provided videos as Supplementary Material. In each of the videos, we show three independent simulations on the aforementioned lattices, with the coupling strengths $(J,K)$ reported in Table S1. Further information about the videos is given below:
  • In the video “MF_lattice.gif” simulations on the mean-field lattice are shown for $N=2000$ spins, where each spin is connected to every other spin. For illustrative purposes, the edges between spins are not shown.
  • In the video “Bethe_lattice.gif” simulations on the Bethe lattice are shown for $N=13121$ spins, corresponding to a Bethe lattice with $8$ layers and a coordination number of $z=4$. For illustrative purposes, only the first $6$ layers of the Bethe lattice are shown.
  • In the video “Square_lattice.gif” simulations on the square lattice are shown for $N=122\times 122$ spins.
S7.3 Eigenvalues of Laplacian matrix